1) Helio Company has two products: A and B. The annual production and sales of Product A is 1,850 units and of Product B is 1,250 units. The company has traditionally used direct labor-hours as the basis for applying all manufacturing overhead to products. Product A requires 0.3 direct labor-hours per unit and Product B requires 0.6 direct labor-hours per unit. The total estimated overhead for next period is $100,485. What is the company’s predetermined overhead rate?

2)

Activities Activity Rates
Assembly $ 14.35 per machine-hour
Processing Orders $ 47.85 per order
inspection $ 70.30 per inspection-hour

2) True Blue Corporation provided the data set forth above from its activity-based costing system. The company makes 430 units of product D28K a year, requiring a total of 690 machine-hours, 40 orders, and 10 inspection-hours per year. The product's direct materials cost is $35.82 per unit and its direct labor cost is $29.56 per unit. What is the unit product cost of product D28K?
Answer #1

1)

Predetermined overhead rate = Estimated overhead / Estimated total labor hours

= 100,485/[(1850*0.3)+(1250*0.6)]

= 100,485/1305

= 77 per DLH

2)

Total Cost

= (35.82*430) + (29.56*430) + (14.35*690) + (47.85*40) + (70.30*10)

= 40,631.90

Unit product cost = 40,631.90/430

= 94.49